OpenStudy (anonymous):

Please help!!! If f(x)=(2x+1)^4, then the fourth derivative of f(x) at x=0 is...

OpenStudy (amistre64):

what is the first derivative?

OpenStudy (anonymous):

There is none that was given.

OpenStudy (amistre64):

im sure there are some rules pertaining to taking a derivative that you can use to find it tho ....

OpenStudy (anonymous):

Do i plug in the 0 to the original function, or to the fourth derivative? that's what I dont know.

OpenStudy (amistre64):

you wait till the fourth; then use x=0 to find the value of f''''(0)

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

youre welcome; i can check your derivatives if need be

OpenStudy (amistre64):

f = (2x+1)^4 f' = 4 (2x+1)^3 2 f'' = 4.3 (2x+1)^2 2.2 f''' = 4.3.2 (2x+1)^1 2.2.2 f'''' = 4.3.2.1 (2x+1)^0 2.2.2.2 x=0 gives f'''' = 4.3.2.1.1.2.2.2.2